The outer faces of a rectangular slab made of equal thickness of iron and brass and maintained at 100degree centrigate and 0 degree respectively. The temperature at the interface is
 

Grade:12th pass

2 Answers

Arun
4 years ago
Rate of flow of heat energy through iron=rate of flow of heat energy through brass
KA(dT)/l= KA(dT)/l
As Area and length are same, therfore,
0.2(100-T)=0.3(T-0)
20-0.2T=0.3T
T=20/0.5
T=40°C
Vikas TU
4 years ago
Dear student 
Please follow the below link to get ans of your question 
Good Luck